ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 191, 2012

Barmpalias, G., J. S. Miller and A. Nies Randomness notions and partial relativization

791

Beiglböck, M. and H. Towsner Transfinite approximation of Hindman’s theorem

41

Bierwirth, H. and M. Siles Molina Lie ideals of graded associative algebras

111

Block, F., A. Gathmann and H. Markwig Psi-floor diagrams and a Caporaso-Harris type recursion

405

Borisenko, A. A. and K. Tenenblat On the total curvature of curves in a Minkowski space

755

Breitner, J. Loop subgroups of F_{r} and the image of their stabilizer subgroups in GL_{r}(ℤ)

545

de la Bretèche, R. and T. D. Browning Binary forms as sums of two squares and Châtelet surfaces

973

Browning, T. D. See R. de la Bretèche Bujalance, E., F.-J. Cirre and P. Turbek Symmetry types of

cyclic covers of the sphere

61

Cirre, F.-J. See E. Bujalance Colombini, F., N. Orrù and L. Pernazza On the regularity of the

roots of hyperbolic polynomials

923

Cordaro, P. D. and N. Hanges Hypoellipticity in spaces of ultradistributions—study of a model case

771

Doré, M. and O. Maleva A compact universal differentiability set with Hausdorff dimension one

889

Downarowicz, T. and Y. Lacroix Forward mean proximal pairs and zero entropy

945

Downey, R. and N. Greenberg Strong jump-traceability II: Ktriviality

647

Evans, M. J. Nielsen equivalence classes of free abelianized extensions of groups

185

Fox, J., C. Lee and B. Sudakov Maximum union-free subfamilies  959

Gathmann, A. See F. Block Gautero, F. and F. Mathéus Poisson boundary of groups acting

on ℝ-trees 585

Giannotti, C. and A. Spiro Steepest descent curves of convex functions on surfaces of constant curvature

279

Gitik, M. and P. Koepke Violating the singular cardinals hypothesis without large cardinals

901

Greenberg, N. See R. Downey Hanges, N. See P. D. Cordaro

Hu, Y. and Y. Shen On quasisymmetric homeomorphisms  209
Huang, X. J. and J. S. Liu Parabolic rectangle packings  667

Kaplan, E. The unramified computation of Rankin–Selberg integrals

for SO_{2l} × GL_{n} 137

Kiessling, J. Properties of cellular classes of chain complexes  483

Koepke, P. See M. Gitik Lacroix, Y. See T. Downarowicz Lee, C. See J. Fox Liu, J. S. See X. J. Huang López Pérez, G. and J. A. Soler Arias The point of continuity

property in Banach spaces not containing ℓ_{1}

347

Maleva, O. See M. Doré Markwig, H. See F. Block Mashevitzky, G. Bases of identities for semigroups of bounded rank

transformations of a set

451

Mathéus, F. See F. Gautero Miller, J. S. See G. Barmpalias Moreno, J. P. and R. Schneider Diametrically complete sets in

Minkowski spaces

701

Neftin, D. Admissibility and field relations  559

Nies, A. See G. Barmpalias Orrù, N. See F. Colombini Paoletti, R. Asymptotics of Szegö kernels under Hamiltonian torus

actions

363

Parshall, B. J. and L. L. Scott Cohomological growth rates and Kazhdan–Lusztig polynomials

85

Pernazza, L. See F. Colombini Plagne, A. Optimally small sumsets in groups IV. Counting multi-

plicities and the λ_{G} functions

739

Pumplün, S. A construction method for some real division algebras with su(3) as derivation algebra

307

Qiu, Y. Normalized local theta correspondence and the duality of inner product formulas 227 Schneider, R. See J. P. Moreno Scott, L. L. See B. J. Parshall

Shelah, S. Many partition relations below density  507
Shelah, S. PCF arithmetic without and with choice  1

Shen, Y. See Y. Hu Siles Molina, M. See H. Bierwirth Soler Arias, J. A. See G. López Pérez Spencer, C. V. and T. D. Wooley Diophantine inequalities and

quasi-algebraically closed fields

721

Spiro, A. See C. Giannotti Sudakov, B. See J. Fox Tenenblat, K. See A. A. Borisenko Towsner, H. See M. Beiglböck Turbek, P. See E. Bujalance

Wise, D. T. Recubulating free groups  337

Wooley, T. D. See C. V. Spencer

Zábrádi, G. Generalized Robba rings  817